Woodstock is currently ten years older than his friend Snoopy. In three years, Woodstock will be three times as old as Snoopy will be. Find their ages.
Answer by richwmiller(17219)

w=s+10
w+3=3(s+3)
s+10+3=3s+9
s+4=3s
4=2s
s=2
w=12
check
15=3*(5)
15=15
ok